D | Question 5 18 pts A charge of +1.7 micro-coulombs lies on a horizontal surface. It has a mass of 3 kg and the coefficient of static friction between it and the surface is 1.6. Another charge of -2.6 micro-coulombs lies on the horizontal surface 8 cm to the right of the charge. A third charge which is negative is located 20 cm above the positive charge on the surface. What is the smallest magnitude this charge needs to be in order for the positive charge to overcome static friction such that the negative charge on the surface begins to move it? Answer in micro-coulombs. 93 qi 9 g2
Answer #1

93 F3 0.20 m F2 10.08 m 1.7 x 10-6 C q2 2.6x 10s c

for the charge to move , the net force on it must be greater than or equal to the static frictional force

Fnet = fs

m = mass of charge = 3 kg

mus = Coefficient of static friction = 1.6

static frictional force on the charge is given as

fs = mus mg   

so

Fnet = fs = mus mg   

Fnet = mus mg   

Fnet =(1.6) (3 x 9.8)

Fnet = 47.04 N

r2 = distance of charge q2 from q1 = 0.08 m

r3 = distance of charge q3 from q1 = 0.20 m

F2 = force by charge q2 on q1 = k q1 q2/r22 = (9 x 109) (1.7 x 10-6) (2.6 x 10-6)/(0.08)2 = 6.22 N

F3 = force by charge q3 on q1 = k q1 q3/r32 = (9 x 109) (1.7 x 10-6) q3/(0.20)2

net force using Pythagorean theorem is given as

Fnet = sqrt(F22 + F23 )

47.04 = sqrt((6.22)2+ F23 )

F3 = 46.6 N

we know that

F3 = (9 x 109) (1.7 x 10-6) q3/(0.20)2

46.6 = (9 x 109) (1.7 x 10-6) q3/(0.20)2

q3 = 121.8 x 10-6 C

magnitude of charge = 121.8 microcoulombs